### v1

# 1CIBM – Center for Biomedical Imaging, Switzerland

Abstract
Quality control (QC) has long been considered essential to guarantee the reliability of neuroimaging studies. It is particularly important for fetal brain MRI, where acquisitions and image processing techniques are less standardized than in adult imaging. In this work, we focus on automated quality control of super-resolution reconstruction (SRR) volumes of fetal brain MRI, an important processing step where multiple stacks of thick 2D slices are registered together and combined to build a single, isotropic and artifact-free T2 weighted volume. We propose , a machine-learning method that extracts more than 100 image quality metrics to predict image quality scores using a random forest model. This approach is well suited to a problem that is high dimensional, with highly heterogeneous data and small datasets. We validate in an out-of-domain (OOD) setting and report high performance (ROC AUC = 0.89), even when faced with data from an unknown site or SRR method. We also investigate failure cases and show that they occur in of the images due to ambiguous configurations for which the rating from the expert is arguable. These results are encouraging and illustrate how a non deep learning-based method like is well suited to this multifaceted problem. Our tool, along with all the code used to generate, train and evaluate the model will be released upon acceptance of the paper.
Keywords:
Quality control Fetal brain MRI Super-resolution reconstruction.1 Introduction
Quality control (QC) seeks to find and discard problematic outputs of a process to prevent problems from propagating [1]. This is fundamental in magnetic resonance imaging (MRI) studies, as insufficient MRI data quality has been shown to bias statistical analyses and neuroradiological interpretation [2]. Image quality control protocols are thus increasingly seen as critical in achieving reliable, generalizable, and replicable results in neuroimaging studies [3, 4]. Automated QC tools are popular for adult brain studies [5, 6, 7]. However, these techniques cannot easily be applied to fetal MRI, as they often rely on priors that are invalid in utero, e.g. assuming that the head is surrounded by air, or rely on models of artifacts that differ from fetal MRI. Indeed, fetal brain MRI typically relies on the acquisition of multiple 2-dimensional (2D) fast spin-echo T2-weighted MR volumes with thick slices [8, 9], followed by a post-processing step known as super-resolution reconstruction (SRR) [10, 11, 12, 13, 14, 15, 16], where stacks of T2w slices are registered into a common space before being combined into a single, high resolution 3D volume (Figure 1 – top left). While these methods include built-in outlier rejection steps to increase their robustness, they can fail, returning a volume that will not be usable for further analysis (Figure 1 – bottom left). These volumes can feature various artifacts, such as topological inconsistencies, geometric distortions, high noise or low contrast between tissues [17] (Figure 1 – right).
Although the quality of SRR volumes greatly influences subsequent processing steps like automatic tissue segmentation and surface extraction, only a few methods have been proposed, and their validation was limited. Closest to this work, the study of [18] used multi-instance deep learning methods for automated quality control on 3D volumes reconstructed using only the approach from [10]. The model trained in this work is not publicly available and the validation was restricted to one acquisition site. Other works have taken a more descriptive approach to the quality of SRR volumes. The work of [19] focused on one popular SRR method called NiftyMIC [14], and proposed to predict the quality of the reconstructed volume based on the percentage of slices from the input stacks that NiftyMIC uses for reconstructing that volume. The work of [17] proposed a protocol for fetal brain MRI quality assessment based on the evaluation of several image and artifact properties, alongside an overall image quality score. However, their sample size remained quite small and they did not perform an automated quality assessment. Finally, while a few automated QC methods have been proposed for low-resolution fetal brain MRI [20, 21, 22], they do not readily apply to SRR volumes, since the artifacts can be quite different.
In this work, we propose to tackle the challenge of automated QC for multi-SRR, multi-site data. Our contributions are as follows: (i) We collect a unique dataset consisting of 673 manual quality ratings combining multiple acquisition sites and multiple SRR methods (ii) We build upon FetMRQC [22] to propose a robust approach based on random forest. We show that this non deep-learning approach, when combined with appropriate data pre-processing, provides a high performance, while requiring only a modest amount of training data. (iii) We evaluate the out-of-domain robustness of our method, focusing on unseen sites and SRR methods. (iv) We study the failure modes of our model and show that 45% of the failures are due to ambiguous ratings, and that nearly all false negatives (wrongly flagged as bad quality) have some level of artifact that, although it is not disturbing to a human rater, is picked up by . The code will be made available upon acceptance of the work. We believe that this method will serve as first open source, easy-to-use baseline for future works in the field.
2 Materials and methods
2.1 Demographics
Data. Fetal brain MRI examinations were retrospectively collected from ongoing research studies at three hospitals: H1 (Lausanne University Hospital, n=141), H2 (BCNatal, n=180) and H3 (University Children’s Hospital Zürich (KISPI), n=37). The study received ethical approval from each center’s institutional review board. These 438 subjects were acquired on different Siemens scanners (Erlangen, Germany) at 1.5T or 3T for H1 and H2, and on different GE Healthcare (Chicago, USA) scanners for H3. As the aim of quality control is to be applicable to a wide range of scenarios, each subject was reconstructed using a randomly assigned set of parameters: a randomized SRR method, target resolution () and selection of input stacks. This resulted in 673 reconstructed volumes using four widely-used SRR methods: SVRTK [11, 23] (n=238), NeSVoR [15] (n=213), NiftyMIC [14] (n=202), MIALSRTK [13] (n=20).
Manual quality ratings. After the SR reconstructions were completed, a manual quality assessment was performed following the protocol of Sanchez et al. [17] (we refer readers to that paper for details on the protocol). The protocol was slightly adapted to focus on categories that were found to have an inter-rater reliability of , leading to the categories illustrated on the right of Figure 1. Global quality rating was obtained as a continuous variable between 0 and 4, and the exclusion threshold was set to 1. Two experts, R1 (20 years of experience in medical image analysis) and R2 (4 years of experience) performed manual quality annotations on all the reconstructions (323 for R1 and 448 for R2, with 98 cases rated twice). The annotation interface allowed to streamline the process, and ratings took a total of 11h40 (7h10 for R1, 4h30 for R2), with a median time of 47 seconds per case. Details on the distribution of SRR quality for the three main methods are provided in Figure 2.
2.2 Automated quality control
Our automated quality control method, that we call , is based on MRIQC [5] and FetMRQC [22]. As illustrated on the right of Figure 1, operates in two steps: 1) extraction of a series of image quality measures (IQMs) from the input volumes, and 2) training of a quality assessment model to predict the reference quality score, based on the IQMs.
Image quality metrics. IQMs are features capturing complementary statistics on a given volume. Well-designed IQMs correlate with various image artifacts, enabling the training of a predictive QC model. However, IQMs that are commonly used in adult MRI are not readily applicable to fetal SRR volumes, as they can rely for instance on air as a background [5], which is inappropriate in utero as the fetus is surrounded by maternal tissues. This prompted us to adapt the IQMs of FetMRQC [22] designed for stacks of 2D slices of T2w fetal brain MRI to the SRR setting. These IQMs are based on image intensity (e.g. mean intensity, SSIM of neighbouring slices, etc.), mask statistics (brain mask centroid coordinates) or segmentation-related statistics (tissue-wise statistics like volume, or contrast between tissue intensities). In addition to these metrics we incorporate new topological IQMs based on Betti numbers and Euler characteristics [24], as this has been shown to be an important metric to quantify the quality of a fetal brain MRI segmentation [25]. Since several IQMs required a brain mask and tissue segmentation, we used the state-of-the-art BOUNTI [26] to compute these maps. In total, 106 IQMs were computed (82 from FetMRQC and 24 from topology). We will provide a detailed description of the IQMs used along with the code.
IQMs preprocessing. Additionally we explore two pre-processing strategies to ensure reliable IQMs computation for all our data. In the first one, we saturate image intensities at 99.5 percentile prior to segmentation, as methods like NeSVoR [15] occasionally produce extremely bright background voxels which in turn leads to a inaccurate segmentation. In the second, we apply brain masking, intensity saturation for low and high values ( and ) and resizing to a 0.8mm resolution prior to the IQM computation.
Predictive model. is based on a random forest, and trained in a binary classification setting, where it predicts whether a data sample has quality below the threshold of , and should be excluded. In addition to it, we considered two variants of . The first one was trained by re-weighted samples according to their underlying continuous label: samples with a quality rating close to 0 (bad quality) or 4 (excellent quality) were given weights of 1, while samples close to the decision threshold (quality score = 1) were given a weight of 0. It is referred to as reweighting. The second one, called Predict_artifacts, made use of the available manual ratings of the type of artifacts present: based on the IQMs, a random forest regressor was trained to predict each specific artifact (cf. Figure 1 right), and then these predictions were used as additional IQMs for the prediction.
2.3 Experiment design
2.3.1 Evaluation setting
Previous studies have shown that domain shift is a major obstacle to the generalization of ML-based methods in medical imaging [27], motivating us to evaluate our models in four different experiments. In-domain (ID), where the data is split by subject across centers; leave-one-site-out cross-validation (CV), where all the data from a single site were used for testing; leave-one-SRR-out cross-validation, where all the data from a given SRR were used for testing; leave-one-site/SRR-out, where all the data from a given site and a given SRR method were left out during training, and the model was tested on the combined data of the left-out site and SRR. This simulates a complex domain shift, where our method is tested on a new site, with data reconstructed using an unseen SRR method. Each experiment was repeated 10 times to account for random variations in the model training and training data sampling.
Metrics. Our experimental setting being a classification task with imbalanced data, we reported the area under the ROC curve (AUC), balanced accuracy (BA), sensitivity, and specificity. Specificity is of particular interest, as we would like to especially avoid false positives, where bad quality data is wrongly predicted as being good.
Baselines. As no public baselines were available, we compared the following models and variants of : (i) Stats. A trivial lower-bound, statistics-based baseline, that predicts everything as good quality or bad quality based on the class frequency. (ii) MLP. A model trained using a multi-layer perceptron (MLP) for prediction instead of a random forest. (iii) reweighting. A variant of using sample re-weighting during training.(iv) pred_artifacts. A more complex variant of , that first predicts the artifacts present in the image and then based on them predicts its quality score. (v) Oracle. An oracle model that uses manual, reference ratings for the specific artifacts for the prediction of the global subjective score, serving as an upper bound on the performance of our method.
3 Results
Ablation study.
We first carried out an in-domain ablation study on the various preprocessing options and variants of . We also considered various thresholds for exclusion from the model’s predicted probabilities, and investigated how changing impacted BA, sensitivity and specificity. Results are presented in Table 1, where we see that using segmentation and IQMs pre-processing reaches the best performance with AUC=0.93 and BA=0.85. Choosing increases the specificity of the model, and we chose to keep it at that level for the rest of the paper. The addition of topological metrics improved the AUC and BA by 3% and 4% respectively. Finally, when looking at the oracle, which uses manual ratings of artifacts as the basis for prediction, we see that our models come close to the predictive ability of the oracle, but fall short by a margin of approximately 7-8%. We also experimented with the site-SRR domain shift removal using ComBaT [28], robust scaling, alternative classifiers like logistic regression, and nested cross-validation but these approaches did not improve the performance and are not reported.
| Model | Prepro. | In-domain generalization | ||||
| Seg. | IQMs | AUC () | BA () | Sens. () | Spec. () | |
| Stats | 0.500.00 | 0.500.00 | 0.360.49 | 0.640.49 | ||
| MLP (t=0.7) | ✓ | ✓ | 0.870.03 | 0.800.04 | 0.850.06 | 0.760.11 |
| (no. topo., t=0.7) | ✗ | ✗ | 0.900.03 | 0.800.04 | 0.760.11 | 0.840.11 |
| (no. topo., t=0.7) | ✓ | ✓ | 0.900.04 | 0.810.04 | 0.780.09 | 0.840.11 |
| (t=0.5) | ✗ | ✗ | 0.910.03 | 0.810.07 | 0.910.06 | 0.710.19 |
| (t=0.5) | ✗ | ✓ | 0.910.03 | 0.800.05 | 0.910.05 | 0.700.13 |
| (t=0.5) | ✓ | ✓ | 0.920.02 | 0.810.06 | 0.910.06 | 0.720.16 |
| (t=0.5) | ✓ | ✓ | 0.930.02 | 0.820.06 | 0.920.05 | 0.720.14 |
| (t=0.7) | ✓ | ✓ | 0.930.02 | 0.850.03 | 0.810.08 | 0.880.07 |
| reweighting (t=0.7) | ✓ | ✓ | 0.930.02 | 0.840.02 | 0.810.08 | 0.880.08 |
| pred_artifacts (t=0.5) | ✓ | ✓ | 0.900.03 | 0.810.04 | 0.860.08 | 0.760.10 |
| Oracle | 0.980.01 | 0.920.02 | 0.890.05 | 0.950.04 | ||
| Model | Prepro. | Site | SRR | Site+SRR | ||||||||||
| Seg. | IQMs | AUC | BA | Sens. | Spec. | AUC | BA | Sens. | Spec. | AUC | BA | Sens. | Spec. | |
| Stats | 0.500.00 | 0.500.00 | 0.360.49 | 0.640.49 | 0.500.00 | 0.500.00 | 0.360.49 | 0.640.49 | 0.500.00 | 0.440.17 | 0.360.49 | 0.530.51 | ||
| MLP (0.7) | ✓ | ✓ | 0.850.09 | 0.790.09 | 0.820.12 | 0.750.24 | 0.860.05 | 0.790.04 | 0.840.09 | 0.730.13 | 0.830.14 | 0.780.10 | 0.820.13 | 0.660.35 |
| ✗ | ✗ | 0.900.05 | 0.760.09 | 0.650.22 | 0.880.13 | 0.860.07 | 0.770.07 | 0.630.13 | 0.900.11 | 0.850.09 | 0.680.11 | 0.530.25 | 0.770.34 | |
| ✗ | ✓ | 0.900.07 | 0.780.09 | 0.680.20 | 0.890.10 | 0.850.06 | 0.760.06 | 0.640.15 | 0.880.10 | 0.840.08 | 0.730.08 | 0.650.22 | 0.740.32 | |
| (t=0.7) | ✓ | ✗ | 0.920.04 | 0.790.09 | 0.690.23 | 0.890.06 | 0.90.04 | 0.80.06 | 0.690.15 | 0.910.07 | 0.890.05 | 0.730.10 | 0.600.27 | 0.770.32 |
| ✓ | ✓ | 0.920.06 | 0.810.09 | 0.740.21 | 0.890.08 | 0.910.06 | 0.790.07 | 0.700.14 | 0.880.09 | 0.880.05 | 0.790.09 | 0.720.22 | 0.750.33 | |
| rew. (t=0.7) | ✓ | ✓ | 0.930.06 | 0.820.09 | 0.750.20 | 0.890.07 | 0.900.06 | 0.790.07 | 0.690.13 | 0.900.08 | 0.890.05 | 0.800.09 | 0.720.23 | 0.770.33 |
| pred_artif. | ✓ | ✓ | 0.900.05 | 0.790.04 | 0.850.14 | 0.730.10 | 0.870.05 | 0.780.05 | 0.860.08 | 0.710.08 | 0.870.07 | 0.750.09 | 0.850.16 | 0.550.26 |
| Oracle | 0.980.02 | 0.920.05 | 0.880.06 | 0.950.05 | 0.970.02 | 0.920.04 | 0.890.08 | 0.940.09 | 0.970.03 | 0.910.05 | 0.890.07 | 0.830.32 | ||
Out-of-domain performance.
In this experiment, we evaluate our model on three OOD settings: leave-one-site-out, -SRR-out and -site/SRR-out. Our results, in Table 2, show an increasing drop in performance across settings: it is less challenging to generalize to an unseen site than to an unseen SRR, and the most challenging setting is the double shift site+SRR. Although the AUC and BA do not vary much for with robust preprocessing, there is a large drop in sensitivity when evaluating the model on a different site (-7%) and SRR (-11%) and a drop in both sensitivity and specificity in the site+SRR case (resp. -9%; -13%). The Site+SRR case also shows a larger variance in sensitivity and specificity of models, suggesting that the model might exhibit some degree of unreliability on this challenging domain shift. Nevertheless, generally retains the best average performance.
Failure modes.
In this experiment, we assess cases where fails. We perform 10 cross-validation experiments with random data splits and identify cases that are consistently mispredicted. We categorized the apparent causes of failure into four groups: bias field, low contrast images, pathological subjects and others.
Out of 673 cases, the prediction was consistently false for 95 (14%)across the repeated experiments. Figure 3 breaks down the distribution of the errors as a function of the manual quality rating. Out of the 95 failed cases, 23 correspond to cases with a strong bias field, and 22 to a low tissue contrast (Figure 3D). Both were challenging for ’s IQMs to capture and accurately classify. This is one limitation of feature engineering approaches that cannot capture all artifacts encountered. This is further illustrated by the cases with streaks and cut brain (Figure 3A and B), that are confidently predicted as good quality, while being excluded by raters. Nevertheless, out of 95 failed cases, 42 had reference quality scores close to the decision boundary and we consider them as ambiguous cases, where a different rater could have assigned a different label (cf. Figure 3C). These cases might be interpreted as limitations from the ground truth data as well as a challenging cases for . Lastly, systematic failure cases also included images of obviously pathological cases, for which the automated segmentation method failed, leading in turn to inaccurate IQMs and mis-classification by the model.
4 Discussion
In this work, we took a look at the complete cycle of data and model in the problem of QC of fetal brain MRI SRR: we did not operate only on improving our model given the data, but also investigated limitations within the data. Along with a rigorous validation of our model in a leave-one-site/SRR-out cross-validation setting, this enabled us to see that our classical, machine learning approach is well suited as a baseline for automated QC for fetal brain MRI SRR. Not only can it perform well even when data from a site and SRR are held out (with an AUC of 0.89), its failure modes can be attributed to the inability of IQMs to capture certain artifacts, as well as to the ambiguity in the data (44% of failures are close to the decision boundary). QC is a multifaceted task, where different causes can lead to a similar rating, and annotations can vary across raters [17], which translates into errors where it is not totally clear whether the problem lies in the prediction or the ground truth score. This variability in the data is further illustrated by the inability of the Oracle baseline to perfectly predict the global quality from specific artifact ratings. Moreover, as depends on the robustness of its components, an inherent limitation originates from the segmentation method used, BOUNTI [23], which is (as expected) unable to segment heavily pathological cases. This is where a deep learning-based approach could be promising: by learning correlations directly from data, it would not require a pre-trained segmentation model, and should capture more diverse artifacts than IQMs. We will investigate this option in future work and will use as a baseline to evaluate the putative gain of a more complex approach.
Overall, this work highlights the importance of considering the entire machine learning development cycle, especially by looking at the data, and at failure cases [29, 30], instead of only focusing on model development. Moreover, while the reported performance metrics are important, they fail to fully capture the behavior of a model, especially on subtle, multifaceted tasks like QC where ratings inherently contain some degree of uncertainty.
4.0.1 Acknowledgements
This work was funded by Era-net NEURON MULTIFACT project (TS: Swiss National Science Foundation grant 31NE30_203977; AM, GA: French National Research Agency, Grant ANR-21-NEU2-0005; EE: Instituto de Salud Carlos III (ISCIII) grant AC21_2/00016 , GMJ: Ministry of Science, Innovation and Universities: MCIN/AEI/10.13039/501100011033/), the SNSF project 215641, and the SulcalGRIDS Project, (GA: French National Research Agency Grant ANR-19-CE45-0014). AJ is supported by the Prof. Max Cloetta Foundation, EMDO Foundation and Vontobel Foundation. TS, VZ and MBC acknowledge access to the facilities and expertise of the CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by CHUV, UNIL, EPFL, UNIGE and HUG.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.